# Entering Real Social World! Benchmarking the Theory of Mind and Socialization Capabilities of LLMs from a First-person Perspective

### v1

@@ 4.1 Data Collection and Validation
The conversion of the third-person perspective ToM benchmark to the first-person perspective is achieved through GPT-4o, followed by manual verification and correction. The game hands for Limit Texas Hold’em and Blackjack card games are generated by RLcard (Zha et al., 2019). Additionally, we manually construct scenarios for both the new world and counterfactual situations. After the data collection, following Chen et al. (2024)’s method, we conduct two rounds of validation to ensure the data’s correctness and quality. In 1st round, author A would first complete all samples created by author B. For stories, questions, and answer options where there are disagreements, authors A and B would discuss and modify them to reach a consensus as much as possible. In 2nd round, for samples where consensus is still not reached, another author C would discuss with authors A and B to determine the final answer. After two rounds of discussion, the final average agreement reaches 97.6%.
4.2 Data Statistics
EgosocialArena includes two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number-Guessing, and Limit Texas Hold’em, totaling 2,195 data entries. A comparison with existing ToM benchmarks is shown in Table 1.
ToM Benchmark Data Volume Evaluation Environment Number of Scenario Perspective
ToMi 999 Static 1 Third-person
HI-ToM 1200 Static 2 Third-person
FanToM 254 Static 2 Third-person
OpenToM 596 Static 3 Third-person
ToMBench 2860 Static 1 Third-person
EgosocialArena (Ours) 2195 Static and Dynamic 7 First-person
Table 1: Comparison of EgosocialArena with existing ToM benchmarks.
@@ 5.1 Experimental Setup
We evaluate a total of 9 popular LLMs, including GPT-4o [^1] https://openai.com/index/hello-gpt-4o/, o1-preview [^2] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620333https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^4] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, LLaMa-3-8B-Instruct-Turbo, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024). To account for the potential influence of model parameters and instruction tuning, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-8B-Instruct-Turbo, as well as LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena. The average accuracy of their responses will serve as the human performance baseline. No extra tutorials or examples are provided to ensure a fair comparison.
@@ 5.2 Evaluation Method
For Daily Life, New World, and CounterFactful scenarios, we present LLMs with a story, a question, and several options, then ask them to pick the correct answer. Using the accuracy of answering questions as the evaluation metric for these scenarios. In the interactive environments of Number Guessing and Texas Hold’em, the evaluation of these scenarios also has standard answers because we propose agents with stable capabilities and behavioral strategies as opponents of LLMs. For the Blackjack scenario, we conducted a manual evaluation. To ensure the quality of the manual evaluation, we measure the average consistency score between evaluators, which reached 96.3%.
Methods ToMI Number Guessing Texas Hold’em
Third-person First-person Level 1 Level 2 Level 3
Open-source Models
LLaMa-3-8B-Chat 50.6 66.2 0.0 0.0 0.0 48.0
LLaMa-3-70B-Chat 58.4 63.2 10.0 20.0 10.0 38.0
LLaMa-3-8B-Instruct 51.1 65.4 0.0 0.0 0.0 44.0
LLaMa-3.1-405B-Instruct 58.0 65.8 80.0 20.0 20.0 56.0
API-based Models
Claude-3-5-Sonnet 71.0 80.5 50.0 10.0 40.0 66.0
GPT-3.5-Turbo 45.5 51.9 10.0 10.0 0.0 56.0
GPT-4-Turbo 55.4 69.7 10.0 20.0 10.0 60.0
GPT-4o 64.1 71.0 10.0 40.0 10.0 62.0
o1-preview 71.9 77.5 90.0 90.0 90.0 72.0
Human
Human Performance 90.2 90.2 90.0 86.0 73.0 82.0
@@
Table 2: The performance of ToM capabilities from the first-person perspective in open-source models and API-based models. The highest and second-highest scores among models and humans in each section are highlighted in blue and red, respectively.
5.3 Experimental Results
The performance of ToM and socialization capabilities from the LLMs’ first-person perspective is shown in Table 2 and 3, respectively. Based on the experimental results, we have obtained some key insights:
@@ 5.3 Experimental Results
Performance Differences in LLMs’ Theory of Mind (ToM) Across Third-Person and First-Person Perspective
@@
As shown in Table 2, all LLMs exhibited improved performance after the ToMI dataset was adapted from a third-person to a first-person perspective. Notably, the Claude and o1-preview models demonstrated significantly stronger ToM capabilities in the first-person perspective compared to other models. Except for GPT-3.5-Turbo, API-based models generally outperformed open-source models, including the recently released LLaMa-3.1-405B-Instruct. However, despite these improvements, there remains a substantial gap between the performance of all LLMs and that of human participants.
Methods New World Counterfact Blackjack ToMBench
All Desire Emotion Intention
Open-source Models
LLaMa-3-8B-Chat 6.7 71.0 84.5 67.4 71.5 73.0 50.0
LLaMa-3-70B-Chat 13.3 59.0 92.5 73.5 72.5 63.5 83.0
LLaMa-3-8B-Instruct 6.7 64.0 81.4 55.1 45.8 70.5 53.5
LLaMa-3.1-405B-Instruct 36.7 66.0 96.5 77.6 75.0 75.5 80.6
API-based Models
Claude-3-5-Sonnet 90.0 74.0 96.9 79.6 72.0 64.5 82.6
GPT-3.5-Turbo 13.3 37.0 89.7 72.5 70.0 86.0 68.5
GPT-4-Turbo 23.3 70.0 95.0 75.6 84.5 67.4 84.0
GPT-4o 36.7 52.0 96.5 85.7 78.0 91.0 96.5
o1-preview 86.7 90.0 97.2 84.8 73.0 91.0 98.0
Human
Human Performance 93.3 91.0 97.0 90.7 83.2 92.0 97.5
Table 3: The performance of socialization capabilities from the first-person perspective of open-source models and API-based models (after the ToMBench conversion, it also includes part of the ToM capability assessment, which we have consolidated here together) with the highest and second-highest scores among models and humans in each section are highlighted in blue and red, respectively.
The powerful capabilities of the o1-preview model are truly surprising
In the Number-Guessing scenario, almost all large language models perform poorly, even in the simplest Level 1 situation, which poses a significant challenge for humans as well. However, the recent o1-preview model has performed exceptionally well, demonstrating a cognitive level in this scenario that approaches or even surpasses that of humans.
The scaling up of open-source models or performing instruction fine-tuning has not yielded significant results
By comparing the performance of LLaMa-3-8B-Chat with LLaMa-3-70B-Chat, as well as LLaMa-3-8B-Chat with LLaMa-3-8B-Instruct models in Table 2, 3, it is observed that neither of these measures has notably improved the ToM and socialization capabilities of LLMs. To enhance LLMs’ ToM and socialization capabilities, innovative approaches will be needed in future research on LLMs.
The performance gap between open-source models and API-based models in socialization scenarios is significant
In counterfactual and New World scenarios, API-based models exhibit extremely flexible cognition, demonstrating a strong ability to adapt to social situations and showing performance comparable to humans. In contrast, open-source models perform poorly. For instance, the LLaMa-3-8B-Chat model only achieves an accuracy rate of 6.7% in New World scenarios, compared to 90.0% for the Claude model.
In the dimensions of desire and intention in social contexts, LLMs have shown potential
As illustrated in Table 3, although LLMs still lag behind humans in the overall dimension of mental states, their performance in the desire and intention dimensions is impressive. These findings can further inspire and serve as a foundation for the application of LLM-based agents in the real social world.
Figure 4: In the scenario of Number Guessing Level 2 (Arithmetic sequence), the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
5.4 Qualitative Analysis
Analysis of Model Failure Causes and Behavioral Patterns in Number Guessing Scenario —— Mid-point Belief, Strange Guess and Get Back on Track
@@ 5.4 Qualitative Analysis
As shown in Figure 4, in the scenario of Number Guessing (Level 2: Arithmetic sequence), we thoroughly investigate the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers. In round 1, with no available information, the GPT-4-Turbo model thinks the opponent will choose the number 50 within the range of 1-100. The same phenomenon is observed in the GPT-3.5-Turbo model, called ”mid-point belief”. Sometimes, the GPT-4-Turbo model continuously believes the opponent will choose progressively smaller numbers throughout the entire interaction, as depicted by the GPT-4-Turbo guess1 curve in Figure 4. Although this is very close to the gold number, it does not capture that the opponent’s chosen numbers form an arithmetic sequence. Another situation occurs when the GPT-4-Turbo model makes a ”strange guess” in the initial rounds, thinking the opponent will suddenly choose larger numbers. After several rounds, it captures that the opponent’s chosen numbers form an arithmetic sequence, called ”Get Back on Track”. Overall, despite the statistical results indicating that the GPT-4-Turbo model does not establish a belief regarding the Level 2 opponent in the Number Guessing scenario, the phenomena we observed suggest that it has started to grasp some patterns.
Analysis of our research and current studies centered on outcomes and benefits
Many current works focus on designing various strategies to improve the performance of LLMs in static or interactive environments. For example, making an LLM display anger to enhance its performance in negotiation scenarios, or using carefully crafted prompt engineering to analyze an opponent’s behavior from multiple perspectives to achieve higher payoffs. Unlike these approaches, we fundamentally explore the ToM capabilities of LLMs from a first-person perspective in various scenarios. Based on the behavior patterns of LLMs that we observed in our experiments, such as the excellent performance of the o1-preview and Claude models in the situated new world scenarios, it indicates that these models have significant potential for further exploration in fields like role-playing and simulation, as well as entering the real social world.
@@ 6 Conclusion
In this paper, considering the limitations of existing ToM and socialization benchmarks, the importance of first-person ToM and socialization capabilities in LLMs, and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and AI agents, we propose the EgoSocialArena framework. This framework is designed to comprehensively evaluate and probe the first-person ToM and socialization capabilities of LLMs in both static and interactive environments, covering multiple scenarios. To avoid the ”babysitting” problem during the evaluation process and achieve fair and comprehensive assessments, we construct rule-based agents at different cognitive levels and train RL agents. We collect a total of 2195 test data entries and test multiple advanced and popular LLMs, revealing some interesting and key insights, and highlighting a significant potential for enhancing the first-person ToM and socialization capabilities of LLMs, allowing them to truly enter the social world.
Appendix A Appendix
A.1 Game Rules
Blackjack
Blackjack, also known as 21, is a popular card game that involves a dealer and a player. Players must decide whether to hit or stand based on their own hand, the dealer’s face-up card, and the dealer’s one hidden card. The objective is to beat the dealer without exceeding 21 points.
Limit Texas Hold’em
The game commences with each player being dealt two private cards, which belong exclusively to the player and remain hidden from the others. Five community cards are then dealt face-up in a series of stages: a three-card Flop, followed by a single card on the Turn and another single card on the River. The player can choose from four actions: Fold, Check, Call, Raise. They aim to construct the best five-card poker hand possible using any combination of their private cards and community cards.
Number Guessing
Each player selects a number between 1 to 100. The objective is to select a number that is closest to 80% of the group’s average choice. The key idea is to guess how others will estimate the average and thus decide the number to submit.
A.2 Limit Texas Hold’em Scenario
Figure 5: Evaluation examples for the Limit Texas Hold’em Scenario.

### v2

@@ 3 Data Collection, Validation and Statistics
The conversion of the third-person perspective to the first-person perspective is achieved through GPT-4o, followed by manual verification and correction. The game hands for Limit Texas Hold’em and Blackjack card games are generated by RLcard (Zha et al., 2019). Defuse bomb environment is based on gym API (Brockman, 2016) and a text interface. Additionally, we manually construct datasets for both the parallel world and counterfactual situations. After the data collection, following Chen et al. (2024b)’s method, we conduct two rounds of validation to ensure the data’s correctness and quality. In 1st round, author A would first complete all samples created by author B. For stories, questions, and answer options where there are disagreements, authors A and B would discuss and modify them to reach a consensus as much as possible. In 2nd round, for samples where consensus is still not reached, another author C would discuss with authors A and B to determine the final answer. After two rounds of discussion, the final average agreement reaches 97.6%. Data statistics of EgoSocialArena are shown in Table 1.
Statistics Number Data Source
Cognitive Intelligence 1235
-Static Cognition 1155 Convertion
-Dynamic Cogntion Evolution-N0.8A 30 Newly Created
-Dynamic Cognition Evolution-Texas 50 Newly Created
Situational Intelligence 675
-Parallel World Situation 90 Newly Created
-Counterfactual Situation 100 Newly Created
-Real Social World Situation 485 Filter, Convertion
Behavioral Intelligence 335
-Adversarial Game 300 Existing
-Cooperative Game 15 Existing
-Social Goal 20 Existing
Table 1: Data Statistics of EgoSocialArena.
@@ 4.1 Experimental Setup
We evaluate a total of eight prominent foundation LLMs, including GPT-4o [^2] https://openai.com/index/hello-gpt-4o/, o1-preview [^3] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620444https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^5] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024). To account for the potential influence of model parameters, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena. The average accuracy of their responses will serve as the human performance baseline. No extra tutorials or examples are provided to ensure a fair comparison. In the behavioral intelligence scenario, we similarly have these students participate in Adversarial Games and Cooperative Games, recording their average performance. For Social-Goal Driven Dialogue scenario, we use the performance of human interactions with GPT-4o as the baseline, given that GPT-4o is the best-performing LLM for this task.
@@ 4.2 Evaluation Method
For the evaluation of static cognition and situational intelligence, we present LLMs with a story, a question, and several options, then ask them to pick the correct answer. Using the accuracy of answering questions as the evaluation metric for these scenarios. For the evaluation of dynamic cognition evolution, these scenarios also has standard answers. For the adversarial and cooperative game scenario, we consider the win rate and team scores. For the Social-goal driven interactive dialogue, we use GPT-4 to automatically evaluate the performance of humans and LLMs in terms of goal completion during their interactions.
Methods Cognitive Intelligence Static Cognition Dynamic Cognition-G0.8A Dynamic Cogntion Third-person First-person $\Delta$ Level 1 Level 2 Level 3 Limit Texas Open-source Models LLaMa-3-8B-Chat 50.6 66.2 +15.6 0.0 0.0 0.0 48.0 LLaMa-3-70B-Chat 58.4 63.2 +4.8 10.0 20.0 10.0 38.0 LLaMa-3.1-405B-Instruct 58.0 65.8 +7.8 80.0 20.0 20.0 56.0 API-based Models Claude-3-5-Sonnet 71.0 80.5 +9.5 50.0 10.0 40.0 66.0 GPT-3.5-Turbo 45.5 51.9 +6.4 10.0 10.0 0.0 56.0 GPT-4-Turbo 55.4 69.7 +14.3 10.0 20.0 10.0 60.0 GPT-4o 64.1 71.0 +6.9 10.0 40.0 10.0 62.0 o1-preview 71.9 77.5 +5.6 90.0 90.0 90.0 72.0 Human Human Performance 90.2 90.2 0.0 90.0 86.0 73.0 82.0
Methods Situational Intelligence Behavioral Intelligence AVG Parallel World Counterfact Real-World Adversarial Cooperative Social Goal Open-source Models LLaMa-3-8B-Chat 6.7 71.0 67.2 51.3 49.7 22.5 34.8 LLaMa-3-70B-Chat 13.3 59.0 73.2 45.0 53.3 25.5 37.3 LLaMa-3.1-405B-Instruct 36.7 66.0 77.3 52.3 65.2 34.0 52.1 API-based Models Claude-3-5-Sonnet 90.0 74.0 79.8 55.0 94.8 50.5 62.8 GPT-3.5-Turbo 13.3 37.0 72.2 46.7 50.3 33.0 34.6 GPT-4-Turbo 23.3 70.0 75.7 54.7 75.6 52.0 47.4 GPT-4o 36.7 52.0 85.8 54.0 80.8 53.0 50.5 o1-preview 86.7 90.0 84.7 56.7 96.3 52.5 80.6 Human Human Performance 93.3 91.0 90.7 56.3 100.0 64.5 91.6
@@
Table 2: Performance of cognitive, situational, and behavioral intelligence from first-person perspective of eight LLMs. Highest and second-highest scores among LLMs and humans in each scenario are highlighted in red and blue, respectively. AVG represents the average value of cognitive, situational, and behavioral intelligence performance.
4.3 Main Results
As shown in Table 2, the o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios. Nevertheless, an 11.0 gap in overall performance remains when compared to the human baseline, leaving plenty of room for model improvement. The second-best performer is the claude-3-5-sonnet model, which demonstrate impressive results in the static cognition and parallel world scenarios. The GPT-4o model performed well in the Real Social World Situation and Social Goal-Driven interactive dialogue scenarios, likely due to being trained with a substantial amount of human feedback. Overall, the performance of open-source models lags significantly behind that of API-based models and most models still exhibit a large performance gap compared to humans. For instance, the LLaMa-3-8B-Chat model achieved an overall score of 34.8, significantly lower than the human performance of 91.6.
4.4 In-Depth Analysis
@@
Performance Differences in LLMs’ ToM Capabilities Across Third-Person and First-Person Perspective
@@
As shown in Table 2, all LLMs exhibited improved performance after the ToMI benchmark is converted from a third-person to a first-person perspective. The Llama3-8B-Chat model achieved the largest improvement of +15.6. Notably, the claude and o1-preview models demonstrated significantly stronger ToM capabilities in the first-person perspective compared to other models. Except for GPT-3.5-Turbo, API-based models generally outperformed open-source models, including the recently released LLaMa-3.1-405B-Instruct. However, despite these improvements, there remains a substantial gap between the performance of all LLMs and human baselines.
The scaling up of open-source models has not yielded significant results
By comparing the performance of LLaMa-3-8B-Chat with LLaMa-3-70B-Chat in Table 2, we observe that although the model size increased significantly, the overall performance on social intelligence improved by only +2.5. We further explore the scaling effects of increasing the size of the LLaMa-3 model on GSM8K (Cobbe et al., 2021) and MMLU (Chung et al., 2024) tasks, finding improvements of +12.9 and +13.4, respectively, as illustrated in Figure 4.
Figure 4: Left: performance evolution corresponding to scaling up LLaMa-3 model size across different task domains. Right: o1-preview model’s output in dynamic cognition evolution—G0.8A scenario.
The powerful mathematical capabilities of the o1-preview model are truly surprising
In the dynamic cognition evolution-G0.8A scenario, almost all LLMs perform poorly, even in the simplest level 1 situation, which poses a significant challenge for humans as well. However, the recent o1-preview model has performed exceptionally well, we analyze its outputs and find that it is highly sensitive to numbers and can capture the correlations between numbers and the underlying patterns behind them, as illustrated in Figure 4. Therefore, when humans are unable to perceive these numerical patterns, the o1-preview model, based on its powerful mathematical capabilities, perceives things that humans have not detected.
Mid-point Belief, Strange Guess and Get Back on Track
@@
As shown in Figure 5, in the scenario of dynamic cognition G0.8A Level 2 (Arithmetic sequence), we thoroughly investigate the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers. In round 1, with no available information, the GPT-4-Turbo model thinks the opponent will choose the number 50 within the range of 1-100. The same phenomenon is observed in the GPT-3.5-Turbo model, called "mid-point belief". Sometimes, the GPT-4-Turbo model continuously believes the opponent will choose progressively smaller numbers throughout the entire interaction, as depicted by the GPT-4-Turbo guess1 curve in Figure 5. Although this is very close to the gold number, it does not capture that the opponent’s chosen numbers form an arithmetic sequence. Another situation occurs when the GPT-4-Turbo model makes a "strange guess" in the initial rounds, thinking the opponent will suddenly choose larger numbers. After several rounds, it captures that the opponent’s chosen numbers form an arithmetic sequence, called Get Back on Track. Overall, despite the statistical results indicating that the GPT-4-Turbo model does not establish a belief regarding the Level 2 opponent in the G0.8A scenario, the phenomena we observed suggest that it has started to grasp some patterns. The belief information for all models across all rounds can be found in Appendix C.
Figure 5: In the scenario of G0.8A Level 2 (Arithmetic sequence), the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
@@ 5 Conclusion
In this paper, considering the social intelligence of LLMs has yet to be systematically organized and studied, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios, incorporating human-machine interaction scenario evaluation is critical and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and LLM-based agents, we propose the EgoToMArena framework. This framework is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, with eight scenarios: static, dynamic cognition; real-world, counterfactual, parallel world situation; cooperative, adversarial game, and social goal-driven human-machine interactive environment, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective. We conduct extensive experiments and observe some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
Limitations
There are three major limitations in our study. (1) Our study only involves the text modality and does not utilize ego-centric images and videos. The social intelligence of Vision-Language Models from a first-person perspective is very important, and we will leave this for future research. (2) Due to the constraint of computing resources and budget, we only evaluate eight prominent foundation LLMs, While we believe that the selected LLMs are representative. (3) Our study evaluates the social intelligence of LLMs from a first-person perspective, a deeper interpretation of these evaluation results from the perspective of explainability research would be more beneficial for the development of LLMs’ social intelligence.
Appendix
Appendix A Related Works
Ego-centric (First-person Perspective) Research
In the fields of computer vision and robotics, there has already been considerable research on a first-person perspective. For example, Cheng et al. (2023) explored whether vision-language models can "Think from a First-person Perspective?" Huang et al. (2023) proposes the construction of embodied agents in a 3D world, which involves acquiring and processing first-person perspective images. Huang et al. (2024) built a bridge between third-person and first-person perspectives at the action level, while Dou et al. (2024) proposed a method designed to transform exocentric video-language data for egocentric video representation learning. However, research on first-person perspectives in the field of natural language processing remains unexplored.
Datasets Related to Social Intelligence
Sap et al. (2022) proposed SocialIQA and used it to evaluate LLMs. SocialIQA contains many questions related to social commonsense. Ziems et al. (2023) introduced NormBank, a large repository of social norms knowledge, which can be used to assess social norm-related tasks. Li et al. (2024) reorganized and classified existing datasets related to social intelligence. Xu et al. (2023) studied LLMs’ understanding of the world and explored how different persuasion strategies could modify LLMs’ worldviews.
Previous evaluations for the ToM of LLMs primarily focus on testing models using narrative stories, also referred to as reading comprehension scenarios. Specifically, Le et al. (2019) proposed the ToMi benchmark based on the classic Sally-Anne test. Wu et al. (2023) introduced the HI-ToM benchmark, which focuses on higher-order belief reasoning and sets up scenarios where agents can communicate with each other. Gandhi et al. (2023) proposed BigToM, which presents a framework for designing a ToM benchmark from synthetic templates for evaluating different aspects of LLMs’ ToM capabilities. Xu et al. (2024) introduced OpenToM, which assigns personalities to agents in the stories and ensures that the storylines are more reasonable and logical. Chen et al. (2024b) proposed ToMBench, which systematically evaluates LLMs across all dimensions of ToM capabilities. Unlike the above methods that require LLMs to read stories and answer related questions, some studies evaluate LLMs’ performance by inputting dialogues to them. Kim et al. (2023) proposed FanToM, which tests LLMs on their ability to infer the mental states of characters in everyday conversations. Chan et al. (2024) introduced NegotiationToM, which restricts the dialogue content to negotiation scenarios.
For the study of LLMs’ behaviors and interaction capabilities, (Agapiou et al., 2022) proposed Melting 2.0, which encompasses various environments such as cooperation and gaming, originally designed for research in multi-agent reinforcement learning. (Zhou et al., 2023) introduced an interactive dialogue environment for large language models under a social goal-driven framework. (Chen et al., 2024a) proposed a game-like environment where different LLMs are paired for competitive interactions.
Strategy Enhancement in Interactive Scenarios
Some work focuses on designing interaction strategies to enable LLMs to gain more benefits during interactions. For example, Zhang et al. (2024b) proposed Agent-pro, Zhang et al. (2024c) introduced K-level reasoning, and Guo et al. (2023) put forward the Suspicion-Agent. Additionally, Li et al. (2023) explored Multi-LLM collaboration by informing LLMs of task rules through prompts. Park et al. (2023) introduced generative agents that can simulate human behavior. Bianchi et al. (2024) explored the social behavior of LLMs in negotiation scenarios. Fu et al. (2023) show LLMs can improve each other in a negotiation scenario. Fan et al. (2024) examined the capability of LLMs to make rational decisions in game theoretic scenarios. Zhang et al. (2024a) propose to optimize the structure of thought.
Necessity of developing LLMs’ Social Intelligence
With LLMs becoming increasingly integrated into our everyday lives, developing LLMs with social intelligence could be better at communicating with us, collaborating with us, understanding us, teaching us and learning from us. (Gandhi et al., 2021, 2023; Rabinowitz et al., 2018; Shu et al., 2021).
Appendix B Case——Limit Texas Hold’em
As illustrated in Figure 6.
Figure 6: A Case for Limit Texas Hold’em.
Appendix C Belief Dynamic Evolution in G0.8A Scenario
The following three tables correspond to the dynamic evolution data of beliefs for various LLMs under Cognitive Levels 3, 2, and 1.
Model Round 1 Round 2 Round 3 Round 4 Round 5 Round 6 Round 7 Round 8 Round 9 Round 10 Accuracy
GPT-4-Turbo 50 45 40 35 30 25 22 17 15 13 0.1
GPT-3.5-Turbo 40 20 60 55 70 90 60 45 75 85 0
GPT-4o 50 40 30 20 15 10 8 6 5 4 0.1
o1-preview 1 20 0.9
Claude-3-5-Sonnet-20240620 65 45 35 28 20 17 14 10 7.5 5.6 0.4
Llama3-8b-chat-hf 67 67 67 67 67 67 67 67 67 67 0
Llama3-70b-chat-hf 50 45 43 30 25 19 15 12 11 7 0.1
Llama3.1-405b-Instruct-Turbo 50 40 35 29 23 19 14.5 11.5 9.5 7.5 0.2
Model Round 1 Round 2 Round 3 Round 4 Round 5 Round 6 Round 7 Round 8 Round 9 Round 10 Accuracy
GPT-4-Turbo 50 45 48 42 36 33 28 22 18 12 0.2
GPT-3.5-Turbo 40 20 60 35 70 50 45 60 45 40 0.1
GPT-4o 50 40 40 30 25 20 15 10 10 5 0.4
o1-preview 1 0.9
Claude-3-5-Sonnet-20240620 65 45 35 25 20 15 12 8 5 8 0.1
Llama3-8b-chat-hf 67 67 67 67 67 67 67 67 67 67 0
Llama3-70b-chat-hf 50 45 38 32 28 24 21 19 16 11 0.1
Llama3.1-405b-Instruct-Turbo 50 40 35 30 28 25 22 18 15 10 0.2
Model Round 1 Round 2 Round 3 Round 4 Round 5 Round 6 Round 7 Round 8 Round 9 Round 10 Accuracy
GPT-4-Turbo 50 45 48 47 48 49 48 47 46 45 0.1
GPT-3.5-Turbo 40 35 70 30 80 40 55 60 50 30 0.1
GPT-4o 50 40 30 40 35 45 45 45 45 45 0.1
o1-preview 1 0.9
Claude-3-5-Sonnet-20240620 65 45 35 25 20 50 50 50 50 50 0.5
Llama3-8b-chat-hf 67 67 67 67 67 67 67 67 67 67 0
Llama3-70b-chat-hf 50 48 52 53 54 55 54 56 57 58 0.1
Llama3.1-405b-Instruct-Turbo 50 33 45 50 50 50 50 50 50 50 0.8